# Feiyu Xiong3, Tong Xu2, Matthew B. Blaschko1

### v1

@@ Abstract
Junyi Zhuthanks: Equal Contribution1, Shuochen Liu2, Yu Yu3, Bo Tangthanks: Correspondence: tangb@iaar.ac.cn3, Yibo Yan4, Zhiyu Li3, Feiyu Xiong3, Tong Xu2, Matthew B. Blaschko1 1KU Leuven, 2 University of Science and Technology of China, 3 Institute for Advanced Algorithms Research, Shanghai, 4 National University of Singapore
@@ 1 Introduction
As illustrated in Figure 1(a), these findings imply that LLMs perform worse when they have high perplexity regarding the context. We also verify that models perform worse in extractive question answering when the perplexity of the answer span in the context is high (see Sec. 3.2). Inspired by these insights, we propose a novel approach named FastMem to improve context awareness via fast memorization of prompts, as depicted in Figure 1(b). Specifically, we adapt an instruction fine-tuned model by maximizing the prompt likelihood before inference. To ensure the efficiency of the optimization and prevent overfitting to small text snippets, we fine-tune only the last FFN module, achieving significant improvements with minimal computational cost. Our contributions are summarized as follows:
1) We propose a novel approach, FastMem, that integrates memorization into the inference process. Specifically, FastMem employs a pretrain-like objective for an instruction fine-tuned model, optimizing it to memorizing the context (i.e., reduce perplexity on the context).
@@ 3.2 Motivation
As introduced in Sec. 1, conflicts between parametric knowledge and contextual information, as well as mass-seeking behavior are supposed to hamper the context awareness. To verify that, we compare the performance of Llama 3-8B-Instruct (Meta AI, 2023) on two datasets: Natural Questions (NQ) (Kwiatkowski et al., 2019) and NQ-SWAP. NQ is constructed based on Wikipedia, which is included in many pre-training corpora. Therefore, the information from NQ has likely been parameterized into Llama 3. NQ-SWAP is constructed from NQ using the method of Longpre et al. (2021), where we swap the answer strings, which are entities, with random similar entities from the corpus. This process creates new context-answer pairs that the model has not observed before.
@@ 4.1 Experimental configurations
Our experiments include evaluations of popular open-source instruction fine-tuned LLMs from the Llama 3 series (8B and 70B) and the Qwen 1.5 series (4B, 7B, and 14B). Additionally, we assess advanced decoding strategies as baseline approaches: Decoding by Contrasting Layers (DoLa) (Chuang et al., 2024) and Context-Aware Decoding (CAD) (Shi et al., 2023).
@@ 4.4 Computation Complexity
Prefix 0.173 46.4 1.029 299.5
NTP 0.058 -- 0.291 --
Table 5: Results of computational complexity. “Memo.” represents the memorization phase of FastMem. “Prefix” refers to the processing of the prefix during inference, and “NTP” stands for next token prediction.
@@ 6 Limitations
Our FastMem approach optimizes the last FFN module, achieving high computational efficiency and performance improvement. Other parameter-efficient approaches, such as LoRA (Hu et al., 2022), may also be applied to extend the optimization to more layers while maintaining high computational efficiency and enhancing performance. This direction was not explored in our work.
@@
When the batch size is set to be larger than one, we optimize the last FFN module to memorize all examples in the batch simultaneously. However, we could also create multiple copies of the last FFN module, with each one adapted to a single example. This direction was not explored in our work.

### v2

@@ Abstract
Junyi Zhuthanks: Equal Contribution1, Shuochen Liu2, Yu Yu3, Bo Tangthanks: Correspondence: tangb@iaar.ac.cn2, Yibo Yan4, Zhiyu Li3, Feiyu Xiong3, Tong Xu2, Matthew B. Blaschko1 1KU Leuven, 2 University of Science and Technology of China, 3 Institute for Advanced Algorithms Research, Shanghai, 4 National University of Singapore
@@ 1 Introduction
As illustrated in Figure 1(a), these findings imply that LLMs perform worse when they have high perplexity regarding the context. We also verify that models perform worse in extractive question answering when the perplexity of the answer span in the context is high (see Sec. 3.2). Inspired by these insights, we propose a novel approach named FastMem to improve context awareness via fast memorization of prompts, as depicted in Figure 1(b). Specifically, we adapt an instruction fine-tuned model by maximizing the prompt likelihood before generation. To ensure the efficiency of the optimization and prevent overfitting to small text snippets, we fine-tune only the last FFN module, achieving significant improvements with minimal computational cost. Our contributions are summarized as follows:
1) We propose a novel approach, FastMem, that integrates memorization into the inference process. Specifically, FastMem employs a pretrain-like objective for an instruction fine-tuned model, optimizing it to memorizing the context provided in the prompt (i.e., reduce perplexity on the context).
@@ 3.2 Motivation
As introduced in Sec. 1, conflicts between parametric knowledge and contextual information, as well as mass-seeking behavior are supposed to hamper the context awareness. To verify that, we compare the performance of Llama 3-8B-Instruct (Meta AI, 2023) on two datasets: Natural Questions (NQ) (Kwiatkowski et al., 2019) and NQ-SWAP. NQ is constructed based on Wikipedia, which is included in many pre-training corpora. Therefore, the information from NQ has likely been parameterized into Llama 3 (Touvron et al., 2023). NQ-SWAP is constructed from NQ using the method of Longpre et al. (2021), where we swap the answer strings, which are entities, with random similar entities from the corpus. This process creates new context-answer pairs that the model has not observed before.
@@ 4.1 Experimental configurations
Our experiments include evaluations of popular open-source instruction fine-tuned LLMs from the Llama 3 series (8B and 70B) and the Qwen 1.5 series (4B, 7B, and 14B). Additionally, we compare our method with advanced decoding strategies: Decoding by Contrasting Layers (DoLa) (Chuang et al., 2024) and Context-Aware Decoding (CAD) (Shi et al., 2023).
@@ 4.4 Computation Complexity
Prefill 0.173 46.4 1.029 299.5
Decode 0.058 -- 0.291 --
Table 5: Results of computational complexity. “Memo.” represents the memorization phase of FastMem. “Prefill” refers to the processing of the prompt during inference, and “Decode” stands for the generation phase.
@@ 6 Limitations
Our FastMem approach optimizes the last FFN module, achieving high computational efficiency and performance improvement. Other parameter-efficient approaches, such as LoRA (Hu et al., 2022), may also be applied to extend the optimization to more layers while maintaining high computational efficiency and enhancing performance. This could be the direction of future work.
@@
When the batch size is set to be larger than one, we optimize the last FFN module to memorize all examples in the batch simultaneously. However, we could also create multiple copies of the last FFN module, with each one adapted to a single example. This could be another possible extension.